A Bibliography of Publications of David M. Young

David M. Young University of Texas at Austin Department of Mathematics

Austin, TX USA

Tel: ?n/a? FAX: ?n/a?

E-mail: young@math.utexas.edu (Internet)

17 November 2023 Version 3.35

Abstract

This bibliography records publications of David M. Young.

Title word cross-reference

100 [YM14]. $6.50 [You53b]. 9 [ALY88]. A [You72a]. A_{1}XB_{1} = C_{1} [NOY01]. A_{2}XB_{2} = C_{2} [NOY01].

-point [ALY88].

13-15 [Cra87a]. 1973 [Gra73, Wat74]. 1979 [VS79, SRS79]. 1983 [ES84, Gar83]. 1984 [Lak84, VS84]. 1986 [D^{+}86]. 1987 [Cra87a, Cra87b]. 1993 [BCEP94]. 1994 [BC95].

2003 [FDKN04]. 2A [KGY80]. 2B [KGRY81]. 2C [KRYG82a, KRYG82b, KORY84]. 2D [KOY89].

3 [Anoxx].

586 [KRYG82a]. 5th [FDKN04].

68 [ACH^{+}68].

A. [You62a]. Abstractions [SDK^{+}95]. Academic [Anoxx, You62a, ACH^{+}68]. Accelerated [HYS76, KGY80, KGRY81, KRYG82a, KRYG82b, You77, You74]. Accelerating [YJK84]. Acceleration [HLY80, MY89, YM14, YJ80, JY88]. accuracy [YM69]. ACM [Cra87a, ACH^{+}68, CHK^{+}65]. Adapting [KY81, KOY82, KY83]. Adaptive [GKMY78, KGMY79, KGY80, KGRY81, KRYG82a, KRYG82b, MY89, JY88]. adjoint [EGRS59a]. Advanced [CKO^{+}84, FDKN04, Gra73, KCO^{+}84]. Advances [VS79, VS84, Anoxx]. Algebra [BD92, You60b, D^{+}86]. algebraic [Rhe72, You72c]. algorithm [Bro87, CYB90, KRYG82a]. Algorithms [Car88, GKMY78, JY86, KGMY79, KY81, KY83, MY87, JY87, YJM86]. Alt [Anoxx]. Alternating [YW63, YW64, YK95, YK96]. alternating-type [YK95, YK96]. Alternative [YC94]. American [Cur56]. Analysis [ALY88, BC95, Bec52, Lak84, Tod62, VWY01, You58, You61b, You73b, Cur56, SSY05]. Angeles [Bec52]. Applications [FDKN04, Gar83, VWY01, Whi76, ISV82, Bec52, NOY01, RSY93, Car88]. Applied [Cur56, HY81, HY04, You97]. Approach [TTY88]. architecture [SDK^{+}95]. Architectures [CKO^{+}84, KCO^{+}84]. Arising [You76]. Arlington [Lak84]. August [BC95, D^{+}86, ES84, FDKN04, Gra73]. Austin [KH90]. Automatic [MY89]. Automonitor [You61a].

Based [TTY88]. Bayesian [SSY05]. Bethlehem [VS79, VS84]. Biconjugate [YC94]. Binomial [YTS89]. Birkhoff [You97]. Blood [TTY88]. Book [Anoxx, You53b, You62a, You93a]. Boston [D^{+}86]. bound [You71]. Boundary [EGRS59a, You74]. brief [KY88]. Bulgaria [BC95, VWY01]. Bureau [Bec52].

C376 [CYB90]. Calculus [TTY88]. Calif [BS84]. California [Bec52]. Carolina [BCEP94]. case [YT01]. CBMS [Rhe72]. CDC [KY83]. Celebrating [Kin04]. Centenary [BCEP94]. Center [Gar83]. Certain [HLY80]. Chebyshev [JY88, MY89, RSY93]. Class [KY75, YK95, YK96]. Colloquium [BC95]. Colorado [SRS79, ISV82]. Combining [CKO^{+}84, KCO^{+}84]. Committee [ACH^{+}68, CHK^{+}65]. common [NOY01]. Comparisons [HTY87]. complementary [SSY05]. Computation [Cra87a, EGRS59a, Ral65, YM69, ???85]. Computational [Ode75]. Computations [KOY84, KOY86a]. Computer [BHK79, CKO^{+}84, KY75, KOY82, KCO^{+}84, SW86, VS79, VS84, You61a, YF62, YF63, ACH^{+}68, CHK^{+}65]. Computers [Anoxx, YHOK85]. Computing [Nas90, NN93, You93a, You93b]. Conditions [JY86, JY87]. Conference [BCEP94, Cra87a, D^{+}86, ES84, FDKN04, Lak84, Rhe72, SW86, VWY01, ???85, ???89, Wat74]. conformal [Bec52, BYZ51, BYZ53, You52]. Congress [Mor70]. Conjugate [JY83, JY86, YJ80, JY87, YJM86]. Conjugate-Gradient [JY83, JY87]. connected [BYZ51]. consistency [You72b]. consistent [You72a]. Construction [Bec52]. Conte [???89]. Control [Gar83, D^{+}86, YM69, You61a]. Convergence [JY53, You69, You70]. Cornelius [BCEP94]. Corporation [Gar83]. covariance [CYB90, YT01]. Cray [Cra87b]. Curriculum [ACH^{+}68, CHK^{+}65, ACH^{+}68]. CYBER [Gar83, KY83].

D [???89]. Dallas [SW86]. Data [Gar83, HTY87, STY87, SSY05, You61a]. David [AK10, Kin04]. December [BCEP94]. dedicated [AK10, RSY93]. Definite [You53a, You72c]. Degree [KY91, YK90, KY92, YK72, You72d, YK93]. Denver [SRS79]. departures [CYB90]. Derived [You75]. Determination [MY89, WY70]. determine [You52]. Developed [KY81]. Development [KY75]. Difference [JY53, KY86, YE59, YW63, You63, YW64, You54]. Differential [CKO^{+}84, KY75, KCO^{+}84, KY84c, KOY86b, VS79, VS84, Wat74, You59, YF62, You63, YF63, YD65a, YD65b, You75, YK84b, Gra73, You50, You53b, You62b, YK85, You93a, You93b]. Diffusion [JY53].

Digital [Ral65]. Direction [YW63, YW64]. Dirichlet [You55]. Discrete [YD65a, YD65b]. Discussion [YK84a, YM69]. displaying [STY87]. distances [Bro87]. Distribution [YTS89, YT01]. domains [BYZ51]. Dr. [KVW10, ???89]. Dual [MY89]. Dundee [Wat74].

Edinburgh [Mor70]. eds [Anoxx]. Effective [BYZ51]. effectiveness [JY88]. Efficient [KY83]. Eigenvalues [EGRS59a]. Element [CKO^{+}84, KCO^{+}84, You76]. Elements [Whi76]. Elliptic [RB85, Sch81, You59, YE59, YF62, YF63, You75, You50, You54, You62b, You74, YXB95, BS84]. ELLPACK [RB85]. Encyclopedia [BHK79]. Engineer [Bec62]. Engineering [Ame64, Cra87b]. ENUMATH [FDKN04]. Equation [JY53, You61b, YW53]. Equations [CKO^{+}84, FEY61, KY75, KCO^{+}84, KY84c, KOY86b, VS79, VS84, Wat74, You59, YE59, YF62, YW63, You63, YF63, YW64, You75, YK84b, Gra73, JY88, NOY01, Rhe72, You50, You54, You62a, You62b, You72c, You73a, YK85, You93a, You93b, YXB95, You53b, You62a]. Equivalence [HLY80]. Error [Ral65, You58]. European [FDKN04]. Exact [You61b]. exceeding [YM14].

Factor [WY70, You71]. Factors [YWD65, YM14]. Fall [SW86]. False [You58]. February [SRS79]. fifth [VS84]. Fifty [Kin04]. Finite [CKO^{+}84, KCO^{+}84, KY86, Whi76, You63, You76]. Fixed [YK71, KY72]. Flight [Gar83]. flows [You52]. FORTRAN [KRYG82a, KGY80, KGRY81, KRYG82b]. Franz [Anoxx]. free [You52]. Frontal [Duf89]. Future [KY84a, KHY94a].

Garrett [You97]. Gene [You93b]. general [NOY01]. Generalizations [CKY99, You72a]. Generalized [JY83, JY86, YJ80, JY87, KCY01]. generated [YXB95]. generating [CYB90]. GMRES [CKY99, KCY01, KYC03]. Goddard [Gar83]. Golub [You93b]. Gradient [JY83, JY86, YJ80, YC94, JY87, YJM86]. graphical [HTY87]. graphs [STY87]. group [YT01]. Guide [GKY78, KORY84, KOY89].

H [You93b]. held [Bec52, Gar83, Gra73, Lak84, VS79, VS84]. Helmholtz [You61b]. High [You88]. Historical [You89, You90, You87]. History [Cra87a, Nas90]. honoring [???89]. Householder [Bro87]. HSNC’87 [Cra87a]. Hydrodynamic [FEY61].

IFIP [ES84, Mor70]. II [BS84, Whi76]. III [VS79]. IMACS [VS79, VS84]. Implementation [KGY80, YK95]. Independence [YT01]. Inequality [YTS89]. INFOMART [SW86]. Information [Mor70]. Institute [Bec52, Gra73]. Interfaces [ES84]. International [BC95, BCEP94, Lak84, VS79, VS84, VWY01, ISV82, Cra87b]. Interpolation [You60a]. introduction [You93a, You93b]. Involving [You61b]. issue [AK10]. Iteration [MY89, You61b, ALY88, You56]. Iterative [BD92, CKO^{+}84, EGRS59a, GKMY78, HLY80, HY81, HY04, KY75, KGMY79, KY79, KGY79, KYG79, KGY80, KY81, KGRY81, KRYG82a, KRYG82b, KY83, KCO^{+}84, KY84c, KOY86b, KCSY88a, KCSY88b, KY91, MY87, You50, You54, YE59, You71, You75, You76, YJ80, YK84b, You88, You89, You90, You03, CKY99, KY92, KCY01, KYC03, YM14, You72b, You72d, YJK84, YK85, You87, YR92, YXB95, YK96, KH90]. ITPACK [GKMY78, GKY78, KGMY79, KGY80, KGRY81, KRYG82a, KOY82, KRYG82b, KY84a, KY84b, KORY85, KY88, KHY94a, KHY94b, YK81, YK84a]. ITPACKV [KORY84, KOY89].

Jacobi [YM14]. James [You93b]. January [SRS79]. Jersey [Cra87a]. Joint [SW86, YT01]. Jr. [AK10, KVW10]. July [Rhe72]. June [Bec52, Lak84, VS79, VS84, VWY01].

Kjeller [Gra73]. Kolmogorov [YTS89].

L [Anoxx, RSY93]. LAN [KYC00]. LAN/MGMRES [KYC00]. Lanczos [BCEP94, JY93]. Lanczos-type [JY93]. LANGMRES [YC94]. Lanham [Gar83]. Laplace [YW53]. Laplacian [ALY88]. Large [HYS76, KGY79, KYG79, KGY80, KGRY81, KRYG82a, KRYG82b, KY83, KCSY88a, KCSY88b, MY87, You71, YK72, You77, YK81, YK90, YC94, KH90, You72d, You72c, YJM86, YR92, YK93, You03]. least [Bro87]. Lecture [Ode75]. Lehigh [VS79, VS84]. Level [You88]. life [KVW10]. Linear [BD92, D^{+}86, HYS76, JY83, KGY79, KYG79, KGY80, KGRY81, KRYG82a, KRYG82b, KY83, KOY84, KOY86a, MY87, You53a, You60b, You71, YK72, You75, You76, You77, YK81, YHOK85, You88, YK90, YK93, YC94, JY88, JY93, KH90, KYC00, Lak84, You56, You72b, You72d, You72c, You73a, YJM86, YR92, You03, D^{+}86].

M [AK10, Kin04, KVW10, You62a, You93b]. mapping [BYZ53, You52]. maps [Bec52]. Maryland [Gar83]. Massachusetts [D^{+}86]. Mathematical [Cur56]. Mathematics [Bec62, Cur56, FDKN04, Mor70, Ode75, Whi76, YG72, YG73, YG88a, YG88b, You97]. Matrices [You53a, CYB90, You72c]. matrix [NOY01]. May [Cra87a]. Mechanics [Ode75]. memory [AK10, RSY93]. Mesh [KY86]. Method [HYS76, KY71, Kin04, WY70, You53a, You58, YB60, You61b, YWD65, YK71, You77, YC94, CKY99, KY72, KCY01, KYC03, YM14, YW53, You71, You74]. Methods [BD92, CKO^{+}84, Car88, EGRS59a, HLY80, HY81, HY04, JY83, KY75, KY79, KGY79, KYG79, KGY80, KGRY81, KRYG82a, KRYG82b, KCO^{+}84, KY84c, KY86, KOY86b, KY91, VS79, VS84, VY89, WY70, YE59, You60b, YF62, YW63, You63, YF63, YW64, You69, YK69, You70, YK72, YJ80, YK84b, You89, You90, YK90, BYZ53, Duf89, HTY87, JY93, KH90, KY92, KYC00, You50, You54, You72b, You72d, YJK84, YK85, You87, YR92, YK93, You93b, YXB95, YK95, YK96, You93a]. MGMRES [KYC00]. Milne [You53b]. Minneapolis [Cra87b]. Minnesota [Cra87b]. misclassification [SSY05]. Modern [Bec62, You73b]. modifications [CKY99]. Modified [YWD65, YK71, KY72]. Modules [ES84, KORY85]. Monterey [BS84]. Morris [Anoxx]. Muller [YB60]. multidimensional [Bro87]. multifrontal [Duf89]. Multilevel [VY89]. multivariate [HTY87, STY87, YT01].

NAA [VWY01]. NASA [Gar83]. National [Bec52]. NATO [Gra73]. Necessary [JY86, JY87]. Non [Lak84]. Non-linear [Lak84]. Nonlinear [Ame64, You75, Rhe72]. nonstationary [YK95]. Nonsymmetric [KY81, YC94, JY93, KYC00]. Nonsymmetrizable [JY83, YJ80, YJK84]. Norms [KY71, YK69]. North [BCEP94]. Norway [Gra73]. Note [YTS89]. Notes [Ode75, You60a, YB60]. November [SW86]. NSF [Rhe72]. NSF-CBMS [Rhe72]. number [YM69]. Numeric [Cra87a]. Numerical [BC95, Bec52, BYZ53, FDKN04, FEY61, Tod62, Wat74, You53b, You59, YE59, You60b, You63, YG72, You73b, YG73, YG88a, YG88b, YW53, You62b, YM69, You93a, You93b, Cur56, FDKN04, Gra73, Rhe72, VWY01].

October [Gar83, KH90]. Omega [YM90]. one [RSY93]. Operators [YD65a, YD65b]. Optimum [WY70, You71]. Order [JY53]. Ordering [KY86, You72a]. ORDVAC [YW53, You55]. Ortega [You93b]. Ostrowski [You62a]. Overrelaxation [KY71, Kin04, WY70, YWD65, You69, YK69, You70, YK71, EGRS59b, KY72, You71]. Overview [You89, KYC00].

P. [RSY93]. PA [Rhe72]. Package [KY75, KGRY81, KRYG82a, KRYG82b, KY84b, YK81, YK84a]. papers [Cra87a, VWY01]. Parabolic [You59, YF62, YF63, You62b]. Parallel [KCSY88a, KCSY88b, VY89, YK95, ???85, YK96, Car88, NN93]. Parallelism [You88]. Parameter [You61b]. Parameters [MY89, YK71, KY72, SSY05]. Partial [CKO^{+}84, KY75, KCO^{+}84, KY84c, KOY86b, VS79, VS84, You59, YF62, YW63, You63, YF63, YW64, YD65a, YD65b, You75, YK84b, Gra73, You50, You54, You62b, YK85]. Past [KY84a, KHY94a]. PDE [ES84, ES84, KGY79, KYG79, KCSY88b]. PDE-Related [KGY79, KYG79]. PDEs [KCSY88a]. Peaceman [You61b]. Pennsylvania [VS79, VS84]. Periodically [YXB95]. Pittsburgh [Rhe72]. Plovdiv [BC95]. point [ALY88]. Points [KY86]. Poisson [SSY05]. Polyhedron [STY87]. polynomials [RSY93]. Position [You58]. Positive [You53a, You72c]. pp [You53b, You62a]. pp. [Anoxx]. Practice [Lak84]. Prague [FDKN04]. Preconditioned [YJM86]. Preface [AK10]. preliminary [CHK^{+}65]. Present [KY84a, KHY94a]. presented [Cra87a]. preserving [YT01]. Press [Anoxx, You62a]. Princeton [Cra87a]. Priori [WY70]. Problem [BS84, Sch81, TTY88, You55]. Problems [Ame64, EGRS59a, RB85, You60b, You61b, You74]. Proc [Mor70]. Procedure [MY89]. Proceedings [Bec52, Cra87b, Cur56, D^{+}86, ES84, FDKN04, Gar83, Gra73, Lak84, SW86, VS79, VS84, ???89, Cra87a, BC95, BCEP94, ISV82, ???85]. Processing [Mor70, ISV82, NN93]. Processors [???85]. program [CHK^{+}65]. programs [ACH^{+}68]. Project [KY84a, KY88]. Properties [You69, You70]. property [You72a]. Purdue [???89].

Rachford [You61b]. Raleigh [BCEP94]. range [YM69]. rate [SSY05]. rational [YR92]. recommendations [CHK^{+}65, ACH^{+}68]. Rectangle [You61b]. Recurrences [KY91, KY92]. Refined [EGRS59a]. Regional [Rhe72]. Related [KGY79, You69, You70, KYG79]. Relaxation [WY70, YWD65, YM14, You71]. Report [GKMY78, ACH^{+}68, CHK^{+}65]. representation [NOY01]. Representations [YD65a, YD65b]. representing [HTY87]. Reservoir [SRS79]. Review [Anoxx, You53b, You62a, You90, You93a, KY88, You87]. Revised [VWY01]. Richardson [You53a, YW53]. Rousse [VWY01]. Routines [KOY82]. Rubinoff [Anoxx].

S [???89]. scaling [Bro87]. scheduled [YM14]. science [ACH^{+}68, CHK^{+}65, ACH^{+}68, CHK^{+}65, Cra87b]. Sciences [BHK79]. Scientific [Cra87a, Nas90, You93b, ???85, NN93, You93a]. Search [You88, YM90]. Second [Bec62, KY91, VWY01, YK72, You72d, YK90, KY92, YK93]. Second-Degree [KY91, YK90, YK72, You72d, KY92, YK93]. Self [EGRS59a]. Self-adjoint [EGRS59a]. Seminar [Gar83]. September [Cra87b]. Series [Bec62]. Several [YWD65, HTY87, RSY93]. signals [D^{+}86]. Simplification [JY83, JY86, JY87]. simply [BYZ51]. Simulation [SRS79]. Single [You61b]. Sixth [Cur56]. smooth [BYZ51]. Society [Cur56]. Söderköping [ES84]. Software [ES84, KY84b, MY87, YK84a, SDK^{+}95, YJM86, Mor70]. Solution [EGRS59a, FEY61, JY53, KORY85, KCSY88a, KCSY88b, MY87, Wat74, You59, You62a, You63, You71, YK72, You73a, You75, You76, YK90, Gra73, NOY01, Rhe72, You53b, YW53, You56, You62b, You72d, You72c, YK93, You03]. Solutions [JY53, You55]. Solve [HYS76]. Solvers [BS84, Sch81, You88]. Solving [KY75, KGY79, KYG79, KGY80, KGRY81, KRYG82a, KRYG82b, KY83, KCO^{+}84, KY84c, You53a, YE59, You60b, YF62, YW63, YF63, YW64, You77, YK84b, YHOK85, YC94, CKO^{+}84, JY88, JY93, KYC00, You50, You54, You74, YK85, YJM86, YR92, YXB95, RB85]. Some [YE59]. SOR [ALY88]. Space [Gar83]. Sparse [GKMY78, KGMY79, KGY79, KYG79, KGY80, KGRY81, KRYG82a, KRYG82b, KY83, KOY84, KOY86a, KCSY88a, KCSY88b, MY87, YK81, YHOK85, You88, YC94, You72c, YJM86, YR92]. Special [AK10]. specified [CYB90]. sphericity [CYB90]. sponsored [Gar83]. squared [Bro87]. squares [Bro87]. SSOR [HYS76, You74, You77]. Standards [Bec52]. State [ISV82]. Stationary [KY91, KY92, YK72, YK90, You72b, YK93]. Storage [GKMY78, KGMY79]. streamlines [You52]. structures [YT01]. Studies [YE59]. Study [Gra73]. subject [SSY05]. Successive [KY71, Kin04, WY70, YWD65, You69, YK69, You70, YK71, KY72, You71]. Sufficient [JY86, JY87]. Supercomputers [KY84c, YK84b, Cra87b, YK85]. Supercomputing [Car88]. support [SDK^{+}95]. surveillance [YM69]. Survey [KY79, Tod62, YF62, YF63, YG72, You73b, YG73, YG88a, YG88b]. Sweden [ES84].

Symmetric

[GKMY78, KGMY79, KY81, You69, You70]. Symposium [Cur56, VS79, VS84, ISV82, Bec52, Cra87b, SRS79]. System [You88]. Systems [ES84, HYS76, JY83, KGY79, KYG79, KGY80, KY81, KGRY81, KRYG82a, KRYG82b, KY83, KOY84, KOY86a, KCSY88a, KCSY88b, MY87, You53a, You61a, You71, YK72, You75, You76, You77, YK81, YHOK85, YK90, YC94, D^{+}86, JY88, JY93, KH90, KYC00, Rhe72, You56, You62a, You72d, You72c, You73a, YJM86, YR92, YK93, You03].

TC [ES84]. Techniques [You76].

Technology [BHK79]. Testing [TTY88]. Texas [Lak84, SW86]. their [RSY93]. them [SDK^{+}95]. Theory [Lak84]. Third [BC95, Cra87b, VS79]. times [KVW10]. tools [SDK^{+}95]. Topics [RSY93]. transformation [BYZ51]. Trends [Lak84, Lak84]. Tutorial [KY86]. two [YT01]. two-group [YT01]. TX [KH90]. type [JY93, You50, You54, YK95, YK96].

undergraduate [CHK^{+}65]. University [Bec52, Lak84, Rhe72, VS79, VS84, ISV82, ???89]. Unsymmetric [You69, You70]. USA [VS79, VS84]. Use [HYS76, KGY79, KYG79, KOY82, KY83, KY84c, YWD65, YK84b, YHOK85, You52, YW53, YK85, YR92]. User [GKY78, KORY84, KOY89]. Using [GKMY78, KGMY79, RB85, You61b, YM14].

V [VS84]. Value [EGRS59a, You74]. variables [RSY93]. Variations [KYC03]. Vector [KOY82, KOY84, KOY86a, KCSY88a, KCSY88b, YHOK85, ISV82, ???85]. Vectorized [KOY86b]. VIth [Lak84]. Vol [Mor70]. Volume [Anoxx, RSY93].

W. [You53b]. Wiley [You53b]. Working [ES84].

Years [Kin04]. Yokohama [NN93]. York [Anoxx, You53b, You62a]. Young [AK10, KVW10, Bro87, EGRS59b, Kin04].

References

xxx:1985:PCV

[???85] ????, editor. Proceedings of the Conference Vector and Parallel Processors for Scientific Computation. Accademia dei Lincei, Roma, Italy, May 1985. ISBN ???? LCCN ????

xxx:1989:PUC

[???89] ????, editor. Purdue University Conference Proceedings honoring Dr. S. D. Conte. ????, ????, 1989. ISBN ???? LCCN ????

Atchison:1968:CRA

[ACH^{+}68] William F. Atchison, Samuel D.

Conte, John W. Hamblen, Thomas E. Hull, Thomas A. Keenan, William B. Kehl, Edward J. McCluskey, Silvio O. Navarro, Werner C. Rheinboldt, Earl J. Schweppe, William Viavant, and David M. Young, Jr. Curriculum 68: Recommendations for academic programs in computer science: a report of the ACM Curriculum Committee on Computer Science. Communications of the Association for Computing Machinery, 11 (3):151–197, March 1968. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Axelsson:2010:PSI

[AK10] Owe Axelsson and David R. Kincaid. Preface [special issue: dedicated to the memory of David M. Young, Jr.]. Numerical linear algebra with applications, 17 (5):741–742, October 2010. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Adams:1988:ASI

[ALY88] Loyce M. Adams, Randall J. LeVeque, and David M. Young. Analysis of the SOR iteration for the 9-point Laplacian. SIAM Journal on Numerical Analysis, 25(5):1156–1180, October 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ames:1964:NPE

[Ame64] William F. Ames, editor. Nonlinear Problems of Engineering. Academic Press, New York, NY, USA, 1964. xiv + 252 pp. LCCN TA331 .A4. Fifteen lectures delivered at a seminar conducted by the Dept. of Mechanical Engineering, University of Delaware, at the university, June 19–21, 1963.

Anonymous:19xx:BRA

[Anoxx] Anonymous. Book review: Advances in Computers, Volume 3. Franz L. Alt and Morris Rubinoff, eds. Academic Press, New York, 1962. 361 pp. Transactions of the American Nuclear Society, ??(??):465, ???? 19xx. CODEN TANSAO. ISSN 0003-

018X. URL https://www.ans.

org/pubs/journals/download/
article-26563/.

Bainov:1995:PTI

[BC95] D. Baĭnov and Valéry Covachev, editors. Proceedings of the Third International Colloquium on Numerical Analysis: Plovdiv, Bulgaria, 13–17 August, 1994. VSP, International Science Publishers, Utrecht, The Netherlands, 1995. ISBN 90-6764-193-6. LCCN QA297 .I45 1994.

Brown:1994:PCL

[BCEP94] J. David Brown, Moody T. Chu, Donald C. Ellison, and Robert J. Plemmons, editors. Proceedings of the Cornelius Lanczos International Centenary Conference, Raleigh, North Carolina, December 12–17, 1993, volume 73 of Proceedings in Applied Mathematics. SIAM Press, Philadelphia, PA, USA, 1994. ISBN 0-89871-339-0. LCCN QC19.2 .C67 1993.

Beauwens:1992:IML

[BD92] R. Beauwens and R. De Groen, editors. Iterative Methods in Linear Algebra. Elsevier Science Publishers (North-Holland), New York, NY, USA, 1992. ISBN 0-444-89248-6. xiii + 636 pp. LCCN QA184 .I44 1991.

Beckenbach:1952:CAC

[Bec52] E. F. Beckenbach, editor. Construction and applications of conformal maps. Proceedings of a symposium held on June 22–25, 1949, at the Institute for Numerical Analysis of the National Bureau of Standards at the University of California, Los Angeles, volume 18 of National Bureau of Standards Applied Mathematics Series. United States Government Printing Office, Washington, DC, USA, 1952. LCCN QA3 .U5 no.18 Bar.

Beckenbach:1962:MME

[Bec62] Edwin Ford Beckenbach, editor. Modern Mathematics for the Engineer: Second Series. McGraw-Hill, New York, NY, USA, 1962. xviii + 456 pp. LCCN QA401 .B4. With an introduction by Magnus R. Hestenes.

Belzer:1979:ECS

[BHK79] Jack Belzer, Albert G. Holzman, and Allen Kent, editors. Encyclopedia of Computer Sciences and Technology. Marcel Dekker, New York, NY, USA, 1979. ISBN 0-8247-2251-5. various pp. LCCN QA76.15 .E5. 32 volumes (1975– 1995).

Browne:1987:YHA

[Bro87] M. W. Browne. The Young– Householder algorithm and the least squares multidimensional scaling of squared distances. Journal of Classification, 4(2): 175–190, 1987. CODEN ???? ISSN 0176-4268 (print), 1432-1343 (electronic).

Birkhoff:1984:EPS

[BS84] Garrett Birkhoff and Arthur Schoenstadt, editors. Elliptic Problem Solvers II (Monterey, Calif., 1983). Academic Press, New York, NY, USA, 1984. ISBN 0-12-100560-7. LCCN QA377 .E533 1983. Proceedings of the Elliptic Problem Solvers Conference, sponsored by the Mathematics and Mechanics Branch, Office of Naval Research, held in Monterey, California, January 10–12, 1983.

Birkhoff:1951:ECT

[BYZ51] G. Birkhoff, D. M. Young, and E. H. Zarantonello. Effective conformal transformation of smooth, simply connected domains. Proceedings of the National Academy of Sciences of the United States of America, 37 (7):411–414, July 1, 1951. CO-DEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

URL www.ncbi.nlm.nih.gov/
pmc/articles/PMC1063389.

Birkhoff:1953:NMC

[BYZ53] G. Birkhoff, D. M. Young, and E. H. Zarantonello. Numerical methods in conformal mapping. In Proceedings of Symposia in Applied Mathematics, Vol. IV, Fluid dynamics, pages 117–140. McGraw-Hill, New York, NY, USA, 1953.

Carey:1988:PSM

[Car88] Graham F. Carey, editor. Parallel Supercomputing: Methods, Algorithms and Applications. Wiley series in parallel computing. John Wiley and Sons, Ltd., New York, London, Sydney, 1988. ISBN 0-471-92436-9. x + 287 pp. LCCN QA76.5 .P314951 1989.

Conte:1965:UPC

[CHK^{+}65] S. D. Conte, John W. Hamblen, Thomas A. Keenan, William B. Kehl, Silvio O. Navarro, Werner C. Rheinboldt, Earl J. Schweppe, David M. Young, Jr., and William F. Atchison. An undergraduate program in computer science — preliminary recommendations. A report from the ACM Curriculum Committee on Computer Science. Communications of the Association for Computing Machinery, 8(9):543–552, September 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Carey:1984:CFE

[CKO^{+}84] Graham F. Carey, David R. Kincaid, Thomas C. Oppe, Kamy Sepehrnoori, and David M. Young. Combining finite element and iterative methods for solving partial differential equations on advanced computer architectures. In Vichnevetsky and Stepleman [VS84], pages 375– 378. LCCN QA377 .I2 1984.

Chen:1999:GMG

[CKY99] Jen-Yuan Chen, David R. Kincaid, and David M. Young. Generalizations and modifications of the GMRES iterative method. Numerical Algorithms, 21(1–4): 119–146, 1999. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Numerical methods for partial differential equations (Marrakech, 1998).

Crane:1987:HAC

[Cra87a] G. E. Crane, editor. HSNC’87:

ACM Conference on the History of Scientific and Numeric Computation, conference proceedings: papers presented at the Conference, Princeton, New Jersey, May 13-15, 1987. ACM Press, New York, NY 10036, USA, October 1987. ISBN 0-89791-229-2. LCCN QA76 .A25 1987.

Cray:1987:SEC

[Cra87b] Cray Research, Inc., editor. Science and engineering on Cray supercomputers: Proceedings of the third international symposium, Minneapolis, Minnesota, September 9–11, 1987. Cray Research, Inc., Minneapolis, MN, 1987. LCCN QA76.5 .I586 1987.

Curtiss:1956:NAP

[Cur56] J. H. Curtiss, editor. Numerical analysis. [Proceedings of the Sixth Symposium in Applied Mathematics of the American Mathematical Society, volume 6 of Proceedings of Symposia in Applied Mathematics. McGraw-Hill, New York, NY, USA, 1956. LCCN QA297 .S9 1953.

Cornell:1990:CAG

[CYB90] J. E. Cornell, D. M. Young, and T. L. Bratcher. C376. An algorithm for generating covariance matrices with specified departures from sphericity. Journal of Statistical Computation and Simulation, 37(3–4):240–243, 1990. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Datta:1986:LAS

[D^{+}86] Biswa Nath Datta et al., editors. Linear algebra in signals, systems, and control: Proceedings of the Conference on Linear Algebra in Signals, Systems, and Control, Boston, Massachusetts, August 12–14, 1986. SIAM Press, Philadelphia, PA, USA, 1986. ISBN 0-89871-223-8. LCCN QA184.C66 1986.

Duff:1989:FMM

[Duf89] Iain S. Duff. Frontal and multifrontal methods. In J. C. Almond and D. M. Young, editors, Proceedings Modern Numerical Algorithms for Supercomputers, pages 445–487. University of Texas at Austin, Austin, TX, USA, 1989.

Engeli:1959:RIM

[EGRS59a] Max Engeli, Theo. Ginsburg, Heinz Rutishauser, and Eduard L. Stiefel. Refined Iterative Methods for Computation of the Solution and the Eigenvalues of Self-adjoint Boundary Value Problems, volume 8. Birkhäuser Verlag, Basel, Switzerland, 1959. 107 pp. LCCN QA3 .Z8 v.8; QA371.

Engeli:1959:YO

[EGRS59b] Max Engeli, Theo. Ginsburg, Heinz Rutishauser, and Eduard L. Stiefel. Young’s overrelaxation. In Refined Iterative Methods for Computation of the Solution and the Eigenvalues of Self-adjoint Boundary Value Problems [EGRS59a], chapter IV.1, page ?? LCCN QA3 .Z8 v.8; QA371.

Engquist:1984:PSM

[ES84] B. Engquist and T. Smedsass, editors. PDE Software: Modules, Interfaces and Systems: Proceedings of the IFIP TC 2 Working Conference on PDE Software– Modules, Interfaces, and Systems, Söderköping, Sweden, 22– 26 August, 1983. Elsevier Science Publishers (North-Holland), New York, NY, USA, 1984. ISBN 0-444-87620-0. LCCN QA377 .I44 1983.

Feistauer:2004:NMA

[FDKN04] M. (Miloslav) Feistauer, Vít Dolejší, Petr Knobloch, and Karel Najzar, editors. Numerical Mathematics and Advanced Applications: Proceedings of ENU-MATH 2003 the 5th European Conference on Numerical Mathematics and Advanced Applications Prague, August 2003. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-642-62288-7, 3-540-21460-7, 3-642-18775-7 (e-book). LCCN QA297 .E89 2003. URL http:/

/link.springer.com/10.1007/
978-3-642-18775-9.

Fyfe:1961:NSH

[FEY61] I. M. Fyfe, R. C. Eng, and D. M. Young. On the numerical solution of the hydrodynamic equations. SIAM Review, 3(4):298–308, ???? 1961. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gary:1983:CAS

[Gar83] J. Patrick Gary, editor. CY-BER 200 Applications Seminar: Proceedings of a seminar sponsored by NASA Goddard Space Flight Center and Control Data Corporation and held in Lanham, Maryland, October 10–12, 1983, volume 2295 of NASA Conference Publication. NASA Scientific and Technical Information Research, Washington, DC, 1983. LCCN QA76.8.C94 C94 1983.

Grimes:1978:IRA

[GKMY78] Roger G. Grimes, David R. Kincaid, William I. MacGregor, and David M. Young. ITPACK report: Adaptive iterative algorithms using symmetric sparse storage. Report CNA-139, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, August 1978.

Grimes:1978:IUG

[GKY78] Roger G. Grimes, David R. Kincaid, and David M. Young. IT-PACK 2.0 user’s guide. Report CNA-150, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, August 1978.

Gram:1973:NSP

[Gra73] J. G. Gram, editor. Numerical solution of partial differential equations: Proceedings of the NATO Advanced Study Institute held at Kjeller, Norway, August 20–24, 1973, volume 2 of NATO advanced study institutes series. Series C: Mathematical and physical sciences. D. Reidel, Dordrecht, Boston, Lancaster, Tokyo, 1973. ISBN 90-277-0413-9. LCCN QA374.N37 1973.

Hageman:1980:ECI

[HLY80] Louis A. Hageman, Franklin T.

Luk, and David M. Young. On the equivalence of certain iterative acceleration methods. SIAM Journal on Numerical Analysis, 17(6):852–873, December 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hamner:1987:CSG

[HTY87] C. G. Hamner, D. W. Turner, and D. M. Young. Comparisons of several graphical methods for representing multivariate data. Computers and Mathematics and Applications, 13(7):647–655, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187901271.

Hageman:1981:AIM

[HY81] Louis A. Hageman and David M.

Young. Applied Iterative Methods. Computer Science and Applied Mathematics, Editor: Werner Rheinboldt. Academic Press, New York, NY, USA, 1981. ISBN 0-12-313340-8. xvii + 386 pp. LCCN QA297.8 .H34 1981.

Hageman:2004:AIM

[HY04] Louis A. Hageman and David M. Young. Applied Iterative Methods. Dover, New York, NY, USA, 2004. ISBN 0-486-43477-X. xx + 386 pp. Unabridged republication of the 1981 original.

Hayes:1976:UAS

[HYS76] L. Hayes, D. M. Young, and E. Schleicher. Use of accelerated SSOR method to solve large linear systems. SIAM Review, 18(4):808, ???? 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

xxx:1982:PSC

[ISV82] Proceedings, International Symposium on Vector Processing Applications (1982: Colorado State University). Colorado State University, Fort Collins, CO, 1982. LCCN ????

Juncosa:1953:OCS

[JY53] M. L. Juncosa and D. M. Young. On the order of convergence of solutions of a difference equation to a solution of the diffusion equation. Journal of the Society for Industrial and Applied Mathematics, 1(2):111–135, December 1953. CODEN JSI-MAV. ISSN 0368-4245 (print), 1095-712X (electronic).

Jea:1983:SGC

[JY83] Kang C. Jea and David M. Young. On the simplification of generalized conjugate-gradient methods for nonsymmetrizable linear systems. Linear Algebra and Its Applications, 52/53:399– 417, 1983.

Joubert:1986:NSC

[JY86] Wayne D. Joubert and David M.

Young. Necessary and sufficient conditions for the simplification of generalized conjugate gradient algorithms. Linear Algebra and Its Applications, 86/87:449–485, 1986.

Joubert:1987:NSC

[JY87] Wayne D. Joubert and David M.

Young. Necessary and sufficient conditions for the simplification of generalized conjugate-gradient algorithms. Linear Algebra and its Applications, 88/89(??):449– 485, 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jea:1988:EAC

[JY88] Kang C. Jea and David M.

Young. On the effectiveness of adaptive Chebyshev acceleration for solving systems of linear equations. Journal of Computational and Applied Mathematics, 24(1–2):33–54, November 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0377042788903421.

Jea:1993:LMS

[JY93] Kang C. Jea and David M.

Young. Lanczos-type methods for solving nonsymmetric linear systems. In Natori and Nodera [NN93], pages 14–24. LCCN ????

Kincaid:1984:CFE

[KCO^{+}84] David R. Kincaid, Graham F. Carey, Thomas C. Oppe, Kamy Sepehrnoori, and David M. Young. Combining finite element and iterative methods for solving partial differential equations on advanced computer architectures. In Vichnevetsky and Stepleman [VS84], pages 375– 378. LCCN QA377 .I2 1984.

Kincaid:1988:VPIa

[KCSY88a] David R. Kincaid, Graham F. Carey, Kamy Sepehrnoori, and David M. Young. Vector and parallel iterative solution of large sparse systems for PDEs. Report CNA-222, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, August 1988.

Kincaid:1988:VPIb

[KCSY88b] David R. Kincaid, Graham F. Carey, Kamy Sepehrnoori, and David M. Young. Vector and parallel iterative solution of large sparse systems for PDE’s. In Cray Research, Inc. [Cra87b], pages 25–44. LCCN QA76.5 .I586 1987.

Kincaid:2001:GGI

[KCY01] David R. Kincaid, Jen-Yuan Chen, and David M. Young. A generalized GMRES iterative method. In Vulkov et al. [VWY01], pages 475–481. ISBN 3-540-45262-1. LCCN ????

Kincaid:1979:IAI

[KGMY79] David R. Kincaid, Roger G.

Grimes, William I. MacGregor, and David M. Young. ITPACK: Adaptive iterative algorithms using symmetric sparse storage. In SRS’79 [SRS79], pages 151–170. ISBN ???? LCCN ????

Kincaid:1981:IFP

[KGRY81] David R. Kincaid, Roger G.

Grimes, John R. Respess, and David M. Young. ITPACK 2B: A Fortran package for solving large sparse linear systems by adaptive accelerated iterative methods. Report CNA-173, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, September 1981. (Also, Report CCSN–44, Computation Center, University of Texas at Austin.).

Kincaid:1979:UIMa

[KGY79] David R. Kincaid, Roger G.

Grimes, and David M. Young. The use of iterative methods for solving large sparse PDE-related linear systems. Mathematics and Computers in Simulation, XXI: 368–375, 1979. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Kincaid:1980:IFI

[KGY80] David R. Kincaid, Roger G.

Grimes, and David M. Young. ITPACK 2A: A Fortran implementation of adaptive accelerated iterative methods for solving large sparse linear systems.

Report CNA-164, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, October 1980.

Kincaid:1990:IML

[KH90] David R. Kincaid and Linda J. Hayes, editors. Iterative methods for large linear systems (Austin, TX, October 19–21, 1988). Academic Press, New York, NY, USA, 1990. ISBN 0-12-407475-8. LCCN QA432 .I84 1990. Papers from a conference held Oct. 19– 21, 1988, at the Center for Numerical Analysis of the University of Texas at Austin. Banquet in honour of David M. Young’s 65th birthday.

Kincaid:1994:IPP

[KHY94a] David R. Kincaid, Linda J. Hayes, and David M. Young. IT-PACK: Past, present, and future. In Colorado Conference on Iterative Methods, volume 1, page ?? University of Colorado & Front Range Scientific Computations, Inc., Boulder, CO, 1994.

Kincaid:1994:ITN

[KHY94b] David R. Kincaid, Linda J. Hayes, and David M. Young. IT-PACK: Then and now. In IMACS 13th World Congress on Computational and Applied Mathematics. IMACS, Department of Computer Science, Rutgers University, New Brunswick, NJ, 1994. To appear.

Kincaid:2004:CFY

[Kin04] David R. Kincaid. Celebrating fifty years of David M. Young’s successive overrelaxation method. In Feistauer et al. [FDKN04], pages 549–558. ISBN 3-642-62288-7, 3-540-21460-7, 3-642-18775-7 (e-book). LCCN QA297 .E89 2003. URL http:/

/link.springer.com/10.1007/
978-3-642-18775-9.

Kincaid:1984:IUG

[KORY84] David R. Kincaid, Thomas C.

Oppe, John R. Respess, and David M. Young. ITPACKV 2C user’s guide. Report CNA-191, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, February 1984.

Kincaid:1985:ISMa

[KORY85] David R. Kincaid, Thomas C.

Oppe, John R. Respess, and David M. Young. ITPACK solution modules. In Rice and Boisvert [RB85], chapter 7, pages 237–258. ISBN 0-387-90910-9, 3-540-90910-9. LCCN QA377 .R53 1985.

Kincaid:1982:AIR

[KOY82] David R. Kincaid, Thomas C.

Oppe, and David M. Young. Adapting ITPACK routines for use on a vector computer. In ISVPA’82 [ISV82], page ?? LCCN ????

Kincaid:1984:VCS

[KOY84] David R. Kincaid, Thomas C.

Oppe, and David M. Young. Vector computations for sparse linear systems. Report CNA-189, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, February 1984.

Kincaid:1986:VCS

[KOY86a] David R. Kincaid, Thomas C. Oppe, and David M. Young. Vector computations for sparse linear systems. SIAM Journal for Discrete and Algebraic Methods, 7(1):99–112, 1986. CO-DEN SJAMDU. ISSN 0196-5212 (print), 2168-345X (electronic).

Kincaid:1986:VIM

[KOY86b] David R. Kincaid, Thomas C. Oppe, and David M. Young. Vectorized iterative methods for partial differential equations. Communications in Applied Numerical Methods, 2(3):289–296, 1986. CODEN CANMER. ISSN 0748-8025 (print), 1555-2047 (electronic).

Kincaid:1989:IUG

[KOY89] David R. Kincaid, Thomas C. Oppe, and David M. Young. IT-PACKV 2D user’s guide. Report CNA-232, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, May 1989.

Kincaid:1982:AIA

[KRYG82a] David R. Kincaid, John R. Respess, David M. Young, and Roger G. Grimes. Algorithm 586: ITPACK 2C: A FOR-TRAN package for solving large sparse linear systems by adaptive accelerated iterative methods. ACM Transactions on Mathematical Software, 8(3):302–322, September 1982. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Kincaid:1982:IFP

[KRYG82b] David R. Kincaid, John R. Respess, David M. Young, and Roger G. Grimes. ITPACK 2C: A Fortran package for solving large sparse linear systems by adaptive accelerated iterative methods. ACM Transactions on Mathematical Software, 8(3):302– 322, 1982. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). Algorithm 586.

Kincaid:2010:LTD

[KVW10] David R. Kincaid, Richard S.

Varga, and Charles H. Warlick. The life and times of Dr. David M. Young, Jr. Numerical linear algebra with applications, 17 (5):743–757, October 2010. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Kincaid:1971:NSO

[KY71] David R. Kincaid and David M.

Young. Norms of the successive overrelaxation method. Report CNA-26, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, July 1971.

Kincaid:1972:MSO

[KY72] David R. Kincaid and David M.

Young. The modified successive overrelaxation method with fixed parameters. Math. Comp., 26 (119):705–717, July 1972. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kincaid:1975:DCP

[KY75] David R. Kincaid and David M. Young. The development of a computer package for solving a class of partial differential equations by iterative methods. Annales de l’Association Internationale pour le Calcul Analogique, 3:186–191, 1975. CODEN AAICAE. ISSN 0020-594X.

Kincaid:1979:SIM

[KY79] David R. Kincaid and David M. Young. Survey of iterative methods. In Belzer et al. [BHK79], pages 354–391. ISBN 0-8247-2251-5. LCCN QA76.15 .E5. 32 volumes (1975–1995).

Kincaid:1981:AIA

[KY81] David R. Kincaid and David M. Young. Adapting iterative algorithms developed for symmetric systems to nonsymmetric systems. In Schultz [Sch81], pages 353–359. ISBN 0-12-632620-7. LCCN QA377 .E53 1980.

Kincaid:1983:AIA

[KY83] David R. Kincaid and David M. Young. Adapting iterative algorithms for solving large sparse linear systems for efficient use on the CDC CYBER 205. In Gary [Gar83], pages 147–160. LCCN QA76.8.C94 C94 1983.

Kincaid:1984:IPP

[KY84a] David R. Kincaid and David M.

Young. The ITPACK project: Past, present, and future. In Birkhoff and Schoenstadt [BS84], pages 53–63. ISBN 0-12-100560-7. LCCN QA377 .E533 1983. Proceedings of the Elliptic Problem Solvers Conference, sponsored by the Mathematics and Mechanics Branch, Office of Naval Research, held in Monterey, California, January 10–12, 1983.

Kincaid:1984:ISP

[KY84b] David R. Kincaid and David M.

Young. The ITPACK software package. In Engquist and Smedsass [ES84], pages 193–206. ISBN 0-444-87620-0. LCCN QA377 .I44 1983.

Kincaid:1984:UIM

[KY84c] David R. Kincaid and David M.

Young. On the use of iterative methods with supercomputers for solving partial differential equations. In Lakshmikantham [Lak84], pages 455–466. ISBN 0-444-87704-5. LCCN QA299.6 .I57 1984.

Kincaid:1986:TFD

[KY86] David R. Kincaid and David M.

Young. A tutorial on finite difference methods and ordering of mesh points. In Stone and Winkler [SW86], pages 556–559. ISBN 0-8186-4743-4 (microfiche), 0-8186-8743-6 (hardcover), 0-8186-0743-2 (paperback). LCCN QA75.5 .F35 1986. IEEE catalog number 86CH2345-7. Computer Society order number 743. ACM order number 401860.

Kincaid:1988:BRI

[KY88] David R. Kincaid and David M. Young. A brief review of the ITPACK project. Journal of Computational and Applied Mathematics, 24(1–2):121– 127, November 1988. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0377042788903470.

Kincaid:1991:SSD

[KY91] David R. Kincaid and David M. Young. Stationary second-degree iterative methods and recurrences. Report CNA-250, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, February 1991.

Kincaid:1992:SSD

[KY92] David R. Kincaid and David M. Young. Stationary seconddegree iterative methods and recurrences. In Beauwens and De Groen [BD92], pages 27– 47. ISBN 0-444-89248-6. LCCN QA184 .I44 1991.

Kincaid:2000:OML

[KYC00] David R. Kincaid, David M. Young, and Jen-Yuan Chen. An overview of MGMRES and LAN/MGMRES methods for solving nonsymmetric linear systems. Taiwanese J. Math., 4(3): 385–396, 2000. ISSN 1027-5487.

Kincaid:2003:VGI

[KYC03] David R. Kincaid, David M.

Young, and Jen-Yuan Chen. Variations of the GMRES iterative method. Applied Numerical Mathematics: Transactions of IMACS, 45(1):3–10, 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). 5th IMACS Conference on Iterative Methods in Scientific Computing (Heraklion, 2001).

Kincaid:1979:UIMb

[KYG79] David R. Kincaid, David M.

Young, and Roger G. Grimes. The use of iterative methods for solving large sparse PDE-related linear systems. In Vichnevetsky and Stepleman [VS79], pages 29– 32. LCCN ????

Lakshmikantham:1984:TTP

[Lak84] V. Lakshmikantham, editor.

Trends in the Theory and Practice of Non-linear Analysis: Proceedings of the VIth International Conference on Trends in the Theory and Practice of Non-Linear Analysis held at the University of Texas at Arlington, June 18–22, 1984. Elsevier Science Publishers (North-Holland), New York, NY, USA, 1984. ISBN 0-444-87704-5. LCCN QA299.6 .I57 1984.

Morrell:1970:IPP

[Mor70] A. J. H. Morrell, editor. Information Processing 68 (Proc. IFIP Congress, Edinburgh, 1968), Vol. 1: Mathematics, Software. North-Holland, Amsterdam, The Netherlands, 1970. ISBN 0-7204-2032-6. LCCN QA 75.5 I57 1968.

Mai:1987:IAS

[MY87] Tsun-Zee Mai and David M. Young. Iterative algorithms and software for the solution of large sparse linear systems. Communications in Applied Numerical Methods, 4(3):435–456, 1987. CODEN CANMER. ISSN 0748-8025 (print), 1555-2047 (electronic).

Mai:1989:DAP

[MY89] Tsun-Zee Mai and David M. Young. A dual adaptive procedure for the automatic determination of iteration parameters for Chebyshev acceleration. International Journal of Numerical Methods in Engineering, 27:483– 499, 1989.

Nash:1990:HSC

[Nas90] Stephen G. Nash, editor. A History of Scientific Computing. ACM Press history series. Addison-Wesley and ACM Press, Addison-Wesley and New York, NY 10036, USA, 1990. ISBN 0-201-50814-1. xix + 359 pp. LCCN QA76.17 .H59 1990.

Natori:1993:PPS

[NN93] Makoto Natori and T. Nodera, editors. Parallel processing for scientific computing (Yokohama, 1993), volume 9 of Advances in numerical methods for large sparse sets of linear equations. Keio University, Yokohama, Japan, 1993. LCCN ????

Navarra:2001:RGC

[NOY01] A. Navarra, P. L. Odell, and D. M. Young. A representation of the general common solution to the matrix equations A_{1}XB_{1} = C_{1} and A_{2}XB_{2} = C_{2} with applications. Computers and Mathematics and Applications, 41(7– 8):929–935, April 2001. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0898122100003308.

Oden:1975:LNM

[Ode75] J. T. Oden, editor. Lecture Notes in Mathematics–Computational Mechanics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. ISBN 0-387-07169-5. vi + 328 pp. LCCN QA3 .L35 v.461.

Rall:1965:EDCb

[Ral65] L. B. Rall, editor. Error in Digital Computation, volume 2. John Wiley and Sons, Ltd., New York, London, Sydney, 1965. Proceedings of an advanced seminar conducted by the Mathematics Research Center, United States Army, at the University of Wisconsin, Madison, October 5–7, 1964.

Rice:1985:SEP

[RB85] John R. Rice and Ronald F.

Boisvert, editors. Solving Elliptic Problems Using ELL-PACK. Springer-Verlag, Berlin, Heidelberg, New York, Tokyo, 1985. ISBN 0-387-90910-9, 3-540-90910-9. x + 497 pp. LCCN QA377 .R53 1985.

Rheinboldt:1972:NSS

[Rhe72] Werner C. Rheinboldt, editor. Numerical solution of systems of nonlinear algebraic equations (NSF-CBMS Regional Conference, University of Pittsburgh, Pittsburgh, PA, July 10–14, 1972). Academic Press, New York, NY, USA, 1972. ISBN 0-12-148950-7. LCCN QA218 .N22 1972.

Rassias:1993:TPO

[RSY93] Themistocle M. Rassias, Hari M. Srivastava, and A. Yanushauskas, editors. Topics in polynomials of one and several variables and their applications: volume dedicated to the memory of P. L. Chebyshev (1821–1894). World Scientific Publishing Co., Singapore, Philadelphia, River Edge, NJ, 1993. ISBN 981-02-0614-3. ix + 638 pp. LCCN QA161.P59 T66 1993. URL http://

                                        www.zentralblatt-math.org/
                                        zmath/en/search/?an=0849.00029.

Schultz:1981:EPS

[Sch81] Martin Schultz, editor. Elliptic Problem Solvers. Academic Press, New York, NY, USA, 1981. ISBN 0-12-632620-7. xiii + 444 pp. LCCN QA377 .E53 1980.

Shaw:1995:ASA

[SDK^{+}95] Mary Shaw, Robert DeLine, Daniel V. Klein, Theodore L. Ross, David M. Young, and Gregory Zelesnik. Abstractions for software architecture and tools to support them. IEEE Transactions on Software Engineering, 21(4):314–335, April 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=385970.

xxx:1979:SRS

[SRS79] Symposium on Reservoir Simulation, Denver, Colorado, January 31, February 1–2, 1979, volume SPE 7687. Society of Petroleum Engineers of AIME, Dallas, TX, 1979. ISBN ???? LCCN ????

Stamey:2005:BAC

[SSY05] J. D. Stamey, J. W. Seaman, and D. M. Young. Bayesian analysis of complementary Poisson rate parameters with data subject to misclassification. Journal of Statistical Planning and Inference, 134(1):36–48, September 1, 2005. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0378375804002058.

Seaman:1987:PGD

[STY87] J. W. Seaman, Jr., D. W. Turner, and D. M. Young. Polyhedron graphs for displaying multivariate data. Computers and Mathematics and Applications, 14(4):269–277, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122187901350.

Stone:1986:FJC

[SW86] H. S. Stone and S. Winkler, editors. Fall Joint Computer Conference Proceedings: November 2–6, 1986, INFOMART, Dallas, Texas. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1986. ISBN 0-8186-4743-4 (microfiche), 0-8186-8743-6 (hardcover), 0-8186-0743-2 (paperback). LCCN QA75.5 .F35 1986. IEEE catalog number 86CH2345-7. Computer Society order number 743. ACM order number 401860.

Todd:1962:SNA

[Tod62]  John Todd, editor.  Survey

of Numerical Analysis. Mc-Graw-Hill, New York, NY, USA, 1962. xvi + 589 pp. LCCN QA297 .T63. URL https:

                                        //babel.hathitrust.org/cgi/
                                        pt?id=mdp.39015000983901.

Turner:1988:CBA

[TTY88] D. W. Turner, F. E. Tidmore, and D. M. Young. A calculus based approach to the blood testing problem. SIAM Review, 30 (1):119–122, March 1988. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Vichnevetsky:1979:ACM

[VS79] R. Vichnevetsky and R. S. Stepleman, editors. Advances in Computer Methods for Partial Differential Equations III: Proceedings of the third IMACS International Symposium on Computer Methods for Partial Differential Equations held at Lehigh University, Bethlehem, Pennsylvania, USA, June 20–22, 1979, volume III. IMACS, Department of Computer Science, Rutgers University, New Brunswick, NJ, 1979. LCCN ????

Vichnevetsky:1984:ACM

[VS84] Robert Vichnevetsky and R. S.

Stepleman, editors. Advances in Computer Methods for Partial Differential Equations V: Proceedings of the fifth IMACS International Symposium on Computer Methods for Partial Differential Equations, held at Lehigh University, Bethlehem, Pennsylvania, USA, June 19–June 21, 1984. IMACS, Department of Computer Science, Rutgers University, New Brunswick, NJ, 1984. LCCN QA377 .I2 1984.

Vulkov:2001:NAA

[VWY01] Lubin Vulkov, Jerzy Waśniewski, and Plamen Yalamov, editors. Numerical Analysis and Its Applications: Second International Conference, NAA 2000, Rousse, Bulgaria, June 11–15, 2000 Revised Papers, volume 1988 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-45262-1. LCCN ????

Vona:1989:PMM

[VY89] Bi Roubolo Vona and David M.

Young. Parallel multilevel methods. In ???? [???89], page ?? ISBN ???? LCCN ????

Watson:1974:CNS

[Wat74] G. A. Watson, editor. Conference on the Numerical Solution of Differential Equations: Dundee 1973, volume 363 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 0-387-06617-9, 3-540-06617-9 (print), 3-540-37914-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3.L28 no.363. URL http:/

                                        /link.springer.com/book/10.
                                        1007/BFb0069119;   http://
                                        www.springerlink.com/content/
                                        978-3-540-37914-0.

Whiteman:1976:MFE

[Whi76] J. R. Whiteman, editor. Mathematics of Finite Elements and Applications II. Academic Press, New York, NY, USA, 1976. ISBN 0-12-747252-5. xiii + 573 pp. LCCN TA347 F5 M37 1975.

Warlick:1970:PMD

[WY70] Charles H. Warlick and David M. Young. A priori methods for the determination of the optimum relaxation factor for the successive overrelaxation method. Report TNN-105, Computation Center, University of Texas at Austin, Austin, TX, USA, May 1970.

Young:1960:NMM

[YB60] David M. Young and Carl Bailey. Notes on Muller’s method. Report TNN-2, Computation Center, University of Texas at Austin, Austin, TX, USA, October 1960.

Young:1994:LAB

[YC94] David M. Young and Jen-Yuan Chen. LANGMRES an alternative to the biconjugate gradient method for solving large sparse nonsymmetric linear systems. In Brown et al. [BCEP94], pages 279–281. ISBN 0-89871-339-0. LCCN QC19.2 .C67 1993.

Young:1965:DRPa

[YD65a] David M. Young and John H.

Dauwalder. Discrete representations of partial differential operators. Report TNN-46, Computation Center, University of Texas at Austin, Austin, TX, USA, May 1965.

Young:1965:DRPb

[YD65b] David M. Young and John H.

Dauwalder. Discrete representations of partial differential operators. In Rall [Ral65], pages 181– 217. Proceedings of an advanced seminar conducted by the Mathematics Research Center, United States Army, at the University of Wisconsin, Madison, October 5– 7, 1964.

Young:1959:SNS

[YE59] David M. Young and Louis Ehrlich. Some numerical studies of iterative methods for solving elliptic difference equations. Report TNN-5, Computation Center, University of Texas at Austin, Austin, TX, USA, April

Young:1962:SCM

[YF62] David M. Young and Thurman G. Frank. A survey of computer methods for solving elliptic and parabolic partial differential equations. Report TNN-20, Computation Center, University of Texas at Austin, Austin, TX, USA, December 1962.

Young:1963:SCM

[YF63] David M. Young and Thurman G. Frank. A survey of computer methods for solving elliptic and parabolic partial differential equations. ICC Bull., 2(??):3– 61, ?? 1963. CODEN ICCBAF. ISSN 0536-1222.

Young:1972:SNM

[YG72] David M. Young and Robert Todd Gregory. A Survey of Numerical Mathematics. Addison-Wesley, Reading, MA, USA, 1972. ISBN 0-201-08773-1, 0-486-65691-8 (Dover paperback). x + 492 (A1–A18 and B1–B14 and I1–I19) pp. LCCN QA297 .Y63 1972.

Young:1973:SNM

[YG73] David M. Young and Robert Todd Gregory. A Survey of Numerical Mathematics. Addison-Wes-ley, Reading, MA, USA, 1973. ISBN 0-486-65692-6 (Dover paperback). 493–1099 + xviii + xvi + xviii pp. LCCN QA297 .Y63 1973.

Young:1988:SNMa

[YG88a] David M. Young and Robert Todd Gregory. A Survey of Numerical Mathematics, volume I. Dover, New York, NY, USA, 1988. ISBN 0-486-65691-8. x + 492 + A22 + B16 + I18 pp. LCCN QA297.Y63 1972. URL http:/

/www.zentralblatt-math.org/
zmath/en/search/?an=0732.65002.

Corrected reprint of the 1973 original.

Young:1988:SNMb

[YG88b] David M. Young and Robert Todd Gregory. A Survey of Numerical Mathematics, volume II. Dover, New York, NY, USA, 1988. ISBN 0-486-65692-6. xii + 607 + A18 + B16 + I18 pp. LCCN QA297.Y63 1972. URL http:/

/www.zentralblatt-math.org/
zmath/en/search/?an=0732.65003.

Corrected reprint of the 1973 original.

Young:1985:UVC

[YHOK85] David M. Young, Linda J. Hayes, Thomas C. Oppe, and David R. Kincaid. On the use of vector computers for solving sparse linear systems. In ???? [???85], page ?? ISBN ???? LCCN ????

Young:1980:GCG

[YJ80] David M. Young and Kang C.

Jea. Generalized conjugate gradient acceleration of nonsymmetrizable iterative methods. Linear Algebra and its Applications, 34(??):159–194, December 1980. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0024379580901652.

Young:1984:ANI

[YJK84] David M. Young, Kang C. Jea, and David R. Kincaid. Accelerating nonsymmetrizable iterative methods. In Birkhoff and Schoenstadt [BS84], pages 323–342. ISBN 0-12-100560-7. LCCN QA377 .E533 1983. Proceedings of the Elliptic Problem Solvers Conference, sponsored by the Mathematics and Mechanics Branch, Office of Naval Research, held in Monterey, California, January 10–12, 1983.

Young:1986:PCG

[YJM86] David M. Young, Kang C. Jea, and Tsun-Zee Mai. Preconditioned conjugate gradient algorithms and software for solving large sparse linear systems. In Datta et al. [D^{+}86], pages 260– 283. ISBN 0-89871-223-8. LCCN QA184.C66 1986.

Young:1969:NSO

[YK69] David M. Young and David R. Kincaid. Norms of the successive overrelaxation methods. Report TNN-94, Computation Center, University of Texas at Austin, Austin, TX, USA, September 1969.

Young:1971:MSO

[YK71] David M. Young and David R. Kincaid. The modified successive overrelaxation method with fixed parameters. Report CNA-33, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, October 1971.

Young:1972:LSS

[YK72] David M. Young and David R.

Kincaid. Linear stationary second-degree methods for the solution of large linear systems. Report CNA-52, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, October 1972.

Young:1981:IPL

[YK81] David M. Young and David R.

Kincaid. The ITPACK package for large sparse linear systems. In Schultz [Sch81], pages 163– 185. ISBN 0-12-632620-7. LCCN QA377 .E53 1980.

Young:1984:ISP

[YK84a] David M. Young and David R.

Kincaid. The ITPACK software package (with a discussion). In Engquist and Smedsass [ES84], pages 193–206. ISBN 0-444-87620-0. LCCN QA377 .I44 1983.

Young:1984:UIM

[YK84b] David M. Young and David R.

Kincaid. On the use of iterative methods with supercomputers for solving partial differential equations. In Lakshmikantham [Lak84], pages 455–466. ISBN 0-444-87704-5. LCCN QA299.6 .I57 1984.

Young:1985:UIM

[YK85] David M. Young and David R.

Kincaid. On the use of iterative methods with supercomputers for solving partial differential equations. In Trends in the theory and practice of nonlinear analysis (Arlington, Tex., 1984), volume 110 of North-Holland Math. Stud., pages 455–466. North-Hol-land, Amsterdam, The Netherlands, 1985.

Young:1990:LSS

[YK90] David M. Young and David R. Kincaid. Linear stationary second-degree methods for solution of large linear systems. Report CNA-244, Center for Numerical Analysis, University of Texas at Austin, Austin, TX, USA, July 1990.

Young:1993:LSS

[YK93] David M. Young and David R. Kincaid. Linear stationary second-degree methods for the solution of large linear systems. In Rassias et al. [RSY93], pages 609–629. ISBN 981-02-0614-3. LCCN QA161.P59 T66 1993. URL http://

www.zentralblatt-math.org/
zmath/en/search/?an=0849.00029.

Young:1995:PIC

[YK95] David M. Young and David R. Kincaid. Parallel implementation of a class of nonstationary alternating-type methods. In Baĭnov and Covachev [BC95], pages 219–222. ISBN 90-6764-193-6. LCCN QA297 .I45 1994.

Young:1996:NCP

[YK96] David M. Young and David R. Kincaid. A new class of parallel alternating-type iterative methods. Journal of Computational and Applied Mathematics, 74(1–2):331–344, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). TICAM Symposium (Austin, TX, 1995).

Young:1969:SCN

[YM69] David M. Young and Alvis E.

McDonald. On the surveillance and control of number range and accuracy in numerical computation (with discussion). In Morrell [Mor70], pages 145–152. ISBN 0-7204-2032-6. LCCN QA 75.5 I57 1968.

Young:1990:SO

[YM90] David M. Young and Tsun-Zee Mai. The search for omega. In Kincaid and Hayes [KH90], pages 293–311. ISBN 0-12-407475-8. LCCN QA432 .I84 1990. Papers from a conference held Oct. 19– 21, 1988, at the Center for Numerical Analysis of the University of Texas at Austin. Banquet in honour of David M. Young’s 65th birthday.

Yang:2014:AJI

[YM14] Xiyang I. A. Yang and Rajat Mittal. Acceleration of the Jacobi iterative method by factors exceeding 100 using scheduled relaxation. Journal of Computational Physics, 274(??):695– 708, October 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0021999114004173.

Young:1950:IMS

[You50] David M. Young. Iterative methods for solving partial differential equations of elliptic type. Ph.D. dissertation, Harvard University, Cambridge, MA, USA, 1950.

Young:1952:UCM

[You52] David M. Young. The use of conformal mapping to determine flows with free streamlines. In Beckenbach [Bec52], pages 125– 136. LCCN QA3 .U5 no.18 Bar.

Young:1953:RMS

[You53a] David Young. On Richardson’s method for solving linear systems with positive definite matrices. Journal of Mathematics and Physics (MIT), 32(1– 4):243–255, April 1953. CO-DEN JMPHA9. ISSN 0097-1421.

                                        URL https://onlinelibrary.
                                        wiley.com/doi/epdf/10.1002/
                                        sapm1953321243.

Young:1953:BRW

[You53b] David M. Young. Book review: W. E. Milne, Numerical solution of differential equations. New York, Wiley, 1953. 11 + 275 pp. $6.50. Bulletin of the American Mathematical Society, 59(6):577–579, ???? 1953. CO-DEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

                                        URL http://projecteuclid.
                                        org/euclid.bams/1183518287.

Young:1954:IMS

[You54] David Young. Iterative methods for solving partial difference equations of elliptic type. Transactions of the American Mathematical Society, 76:92–111, 1954. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (electronic).

Young:1955:OSD

[You55] David M. Young. ORDVAC solutions of the Dirichlet problem. Journal of the ACM, 2(3):137– 161, July 1955. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Young:1956:SLS

[You56] David M. Young. On the solution of linear systems by iteration. In Curtiss [Cur56], pages 283–298. LCCN QA297 .S9 1953.

Young:1958:EAM

[You58] David M. Young. Error analysis for the method of false position. Report TNN-1, Computation Center, University of Texas at Austin, Austin, TX, USA, October 1958.

Young:1959:NSE

[You59] David M. Young. The numerical solution of elliptic and parabolic partial differential equations. Report TNN-4, Computation Center, University of Texas at Austin, Austin, TX, USA, April 1959.

Young:1960:NI

[You60a] David M. Young. Notes on interpolation. Report TNN-3, Computation Center, University of Texas at Austin, Austin, TX, USA, November 1960.

Young:1960:NMS

[You60b] David M. Young. Numerical methods for solving problems in linear algebra. Report TNN-9, Computation Center, University of Texas at Austin, Austin, TX, USA, May 1960.

Young:1961:ASC

[You61a] David M. Young. Automonitor systems for the Control Data 1604 computer. Report TNN-14, Computation Center, University of Texas at Austin, Austin, TX, USA, August 1961.

Young:1961:EAP

[You61b] David M. Young. Exact analysis of the Peaceman–Rachford method using a single iteration parameter for problems involving the Helmholtz equation in the rectangle. Report TNN-12, Computation Center, University of Texas at Austin, Austin, TX, USA, June 1961.

Young:1962:BRM

[You62a] David M. Young. Book review: A. M. Ostrowski, Solution of equations and systems of equations. Academic Press, New York, 1960, 9 + 202 pp. Bulletin of the American Mathematical Society, 68(4):306– 308, 1962. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic). URL http:

                                        //projecteuclid.org/euclid.
                                        bams/1183524656.

Young:1962:NSE

[You62b] David M. Young, Jr. The numerical solution of elliptic and parabolic partial differential equations. In Todd [Tod62], chapter 11, pages 380–438. LCCN QA297 .T63. URL https:

//babel.hathitrust.org/cgi/
pt?id=mdp.39015000983901.

Young:1963:NSP

[You63] David M. Young. On the numerical solution of partial differential equations by finite difference methods. Report TNN-21, Computation Center, University of Texas at Austin, Austin, TX, USA, January 1963.

Young:1969:CPS

[You69] David M. Young. Convergence properties of the symmetric and unsymmetric successive overrelaxation methods and related methods. Report TNN-96, Computation Center, University of Texas at Austin, Austin, TX, USA, September 1969.

Young:1970:CPS

[You70] David M. Young. Convergence properties of the symmetric and unsymmetric successive overrelaxation methods and related methods. Math. Comp., 24(112): 793–807, October 1970. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Young:1971:ISL

[You71]  David M. Young.  Iterative

Solution of Large Linear Systems. Computer Science and Applied Mathematics, Editor: Werner Rheinboldt. Academic Press, New York, NY, USA, 1971. ISBN 0-12-773050-8. xxiv + 570 pp. LCCN QA195 .Y681 1971.

Young:1972:GPC

[You72a] David M. Young. Generalizations of property A and consistent ordering. SIAM Journal on Numerical Analysis, 9(3):454–463, September 1972. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Young:1972:CLS

[You72b] David M. Young. On the consistency of linear stationary iterative methods. SIAM Journal on Numerical Analysis, 9(1):89– 96, March 1972. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Young:1972:SLS

[You72c] David M. Young. On the solution of large systems of linear algebraic equations with sparse, positive definite matrices. In Rheinboldt [Rhe72], pages 101– 156. ISBN 0-12-148950-7. LCCN QA218 .N22 1972.

Young:1972:SIM

[You72d] David M. Young. Second-degree iterative methods for the solution of large linear systems. Journal of Approximation Theory, 5(2):137– 148, March 1972. CODEN JAX-TAZ. ISSN 0021-9045 (print), 1096-0430 (electronic). Collection of articles dedicated to J. L. Walsh on his 75th birthday, II (Proc. Internat. Conf. Approximation Theory, Related Topics and Appl., Univ. Maryland, College Park, Md., 1970).

Young:1973:SLS

[You73a] David M. Young. Solution of linear systems of equations. In Gram [Gra73], pages 35– 54. ISBN 90-277-0413-9. LCCN QA374.N37 1973.

Young:1973:SMN

[You73b] David M. Young. A survey of modern numerical analysis. SIAM Review, 15(2):503–523, April 1973. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Young:1974:ASM

[You74] David M. Young. On the accelerated SSOR method for solving elliptic boundary value problems. Lecture Notes in Mathematics, 363:176, 1974. CODEN LNMAA2. ISBN 3-540-06617-9 (print), 3-540-37914-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/
10.1007/BFb0069135.

Young:1975:ISL

[You75] David M. Young. Iterative solution of linear and nonlinear systems derived from elliptic partial differential equations. In Oden [Ode75], pages 265–296. ISBN 0-387-07169-5. LCCN QA3 .L35 v.461.

Young:1976:ISL

[You76] David M. Young. Iterative solution of linear systems arising from finite element techniques. In Whiteman [Whi76], pages 439– 464. ISBN 0-12-747252-5. LCCN TA347 F5 M37 1975.

Young:1977:ASM

[You77] David M. Young. On the accelerated SSOR method for solving large linear systems. Advances in Mathematics, 23(3):215–271, 1977. CODEN ADMTA4. ISSN 0001-8708 (print), 1090-2082 (electronic).

Young:1987:HRI

[You87]  D. M. Young.  An histori-

cal review of iterative methods. In Crane [Cra87a], pages 117– 123. ISBN 0-89791-229-2. LCCN QA76 .A25 1987.

Young:1988:SHL

[You88] David M. Young. The search for ‘high level’ parallelism for iterative sparse linear system solvers. In Carey [Car88], chapter 7, page ?? ISBN 0-471-92436-9. LCCN QA76.5 .P314951 1989.

Young:1989:HOI

[You89] David M. Young. A historical overview of iterative methods. Computer Physics Communications, 53(1–3):1–17, May 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0010465589901458.

Young:1990:HRI

[You90] David M. Young. A historical review of iterative methods. In Nash [Nas90], pages 180– 194. ISBN 0-201-50814-1. LCCN QA76.17 .H59 1990.

Young:1993:BRS

[You93a] David M. Young. Book review:

Scientific computing and differential equations—An introduction to numerical methods. Bulletin of the American Mathematical Society (new series), 28(2):397– 398, 1993. CODEN BAMOAD. ISSN 0273-0979 (print), 1088-9485 (electronic).

Young:1993:SCD

[You93b] David M. Young. Scientific computing and differential equations — an introduction to numerical methods by Gene H. Golub and James M. Ortega. Bulletin (new series) of the American Mathematical Society, 28(2):397, April 1993. ISSN 0273-0979 (print), 1088-9485 (electronic).

Young:1997:GBA

[You97]  David M. Young.  Garrett

Birkhoff and applied mathematics. Notices of the American Mathematical Society, 44 (11):1446–1450, 1997. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Young:2003:ISL

[You03] David M. Young. Iterative solution of large linear systems. Dover, New York, NY, USA, 2003. ISBN 0-486-42548-7. xxvi + 570 pp. Unabridged republication of the 1971 edition [Academic Press, New York-London, MR 305568].

Young:1971:BOR

[You71] David M. Young. A bound for the optimum relaxation factor for the successive overrelaxation method. Numer. Math., 16 (??):408–413, ?? 1970/71. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Young:1992:URI

[YR92] David M. Young and Bi Roubolo Vona. On the use of rational iterative methods for solving large sparse linear systems. Applied Numerical Mathematics: Transactions of IMACS, 10(3– 4):261–278, ?? 1992. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). A Festschrift to honor Professor Garrett Birkhoff on his eightieth birthday.

Young:2001:IDP

[YT01] D. M. Young and D. W. Turner. Independence distribution preserving joint covariance structures for the multivariate twogroup case. Statistics: a Journal of Theoretical and Applied Statistics, 35(4):331–347, 2001. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Young:1989:NKI

[YTS89] D. M. Young, D. W. Turner, and J. W. Seaman. A note on a Kolmogorov inequality for the binomial distribution. Theory of Probability and its Applications, 33(4):747–749, ???? 1989. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Young:1953:URM

[YW53] David M. Young and C. H. Warlick. On the use of Richardson’s method for the numerical solution of Laplace’s equation on the ORDVAC. Memorandum Report 709, Ballistic Research Laboratories, Aberdeen Proving Ground, MD, USA, 1953. ?? pp.

Young:1963:ADM

[YW63] David M. Young and Mary Fanett Wheeler. Alternating direction methods for solving partial difference equations. Report TNN-30, Computation Center, University of Texas at Austin, Austin, TX, USA, December 1963.

Young:1964:ADM

[YW64] David M. Young and Mary Fanett Wheeler. Alternating direction methods for solving partial difference equations. In Ames [Ame64], pages 220–246. LCCN TA331 .A4. Fifteen lectures delivered at a seminar conducted by the Dept. of Mechanical Engineering, University of Delaware, at the university, June 19–21,

Young:1965:UMS

[YWD65] David M. Young, Mary F. Wheeler, and James A. Downing. On the use of the modified successive overrelaxation method with several relaxation factors. Report TNN-39, Computation Center, University of Texas at Austin, Austin, TX, USA, January 1965.

Young:1995:PGI

[YXB95] David M. Young, Shengyou Xiao, and Karen R. Baker. Periodically generated iterative methods for solving elliptic equations. Applied Numerical Mathematics: Transactions of IMACS, 19(3): 375–387, 1995. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Special issue on iterative methods for linear equations (Atlanta, GA, 1994).